PDiddi:

Hi! So, I have a question regarding a geometry problem around Special Right Triangles (45,45,90 and 30,90,60 triangles) the problem also requires the Pythagorean theorem: A farmer’s conveyor belt carries bales of hay from the ground to the barn loft. The conveyor belt moves at 100 ft/min. How many seconds does it take for a bale of hay to go from the ground to the loft? The triangle is 30 90 60 one and the shortest leg is 12 ft.

surjithayer:

shortest leg in 30,90,60 triangle=12 ft hypotenuse=2*shortest leg=2*12=24ft time taken to cover 24 ft\[=\frac{ 1 }{ 100 }\times24=\frac{ 24 }{ 100 }~\min=\frac{ 24 }{ 100 }\times60=\frac{ 144 }{ 10 }=14.4 ~\sec\]

PDiddi:

I got the same answer too, but it says to answer in seconds, which is around 0.24...does that make sense?

PDiddi:

oh ok nvm thx!

surjithayer:

no,it is 14.4 sec.

PDiddi:

Yes...thx!